Shamrock, Inc. is considering these two alternatives to finance its construction of a new $1.65 million plant:
1. Issuance of 165,000 shares of common stock at the market price of $10 per share.
2. Issuance of $1.65 million, 6% bonds at face value.
Complete the table. (Round earnings per share to 2 decimal places, e.g. $2.66.)

Issue Stock

Issue Bonds

Income before interest and taxes

$1,595,000 $1,595,000

Interest expense from bonds

enter a dollar amount

enter a dollar amount

Income before income taxes

enter a subtotal of the two previous amounts

enter a subtotal of the two previous amounts

Income tax expense (40%)

enter a dollar amount

enter a dollar amount

Net income

$enter a total net income

$enter a total net income

Outstanding shares

enter a number of shares

660,000

Earnings per share

$enter earnings per share rounded to 2 decimal places

$enter earnings per share rounded to 2 decimal places

LINK TO TEXT

Indicate which alternative is preferable.

select between 2 options

Issuance of stockIssuance of bonds

is preferable.
Answer #1

Answer-

SHAMROCK INC.
Effect on net income $ earnings per share
Particulars Plan A Plan B
Issue Stock Issue Bonds
$ $
Income before interest & taxes 1595000 1595000
Less:- Interest on bonds $1650000*6% 0 99000
Income before taxes 1595000 1496000
Income Taxes 40% $1595000*40%=$638000 $1496000*40% =$598400
Income after taxes(A) 957000 897600
No of shares (B) 660000 Shares+165000 Shares (New issue)=825000 660000
Earning per share C=A/B 1.16 1.36

Issuance of bonds would be preferable due to higher earning per share compare with the issue of shares.